OpenStudy (anonymous):

Simplify 15 - 12 - 8 + 4 - 6 + 1 is the answer -4??

OpenStudy (anonymous):

i got -6

OpenStudy (anonymous):

(15 - 12) - 8 + 4 - 6 + 1= (3 - 8) + 4 - 6 + 1= (-5 + 4) - 6 + 1= (-1 - 6) + 1 = -7 + 1 = -6